# Approximating the coefficients of the Bessel functions

### v3

@@ 10.4. The uniform convergence of Bessel functions for Vershik-Kerov
In this subsection, we reprove the results of the papers [assiotis, rank_infinity] and extend these results to the $D^{N}$ root system. The setting that we consider is equivalent to the setting of Vershik-Kerov sequences that the papers consider, see Remark˜10.14. First, we prove the analogues of Theorems˜6.10, 7.4 and 8.21.
@@
Suppose $F_{N}(x_{1},\ldots,x_{N})=\exp\left(\sum_{\lambda\in\Gamma_{\text{even};\,N}}c_{\lambda}(N)p_{\lambda}\right)+e\exp\\ \left(\sum_{\lambda\in\Gamma_{\text{even};\,N}}d_{\lambda}(N)p_{\lambda}\right)$ for all $N\geq 1$. Assume that $\theta\in\mathbb{C}$ is fixed. Consider the following statements.
@@
    For all $\lambda\in\Gamma_{\text{even}}$, $\lim_{N\rightarrow\infty}d_{\lambda}(N)=d_{\lambda}\in\mathbb{C}$.
@@
    For all $\nu\in\Gamma_{\text{even}}$
@@
Theorem˜10.13 is an example of an application of the results of this paper. Part (A) of the theorem appears in [assiotis] while parts (A) and (B) appear in [rank_infinity]. Part (C) has not appeared previously, although it is straightforward to deduce. See Remark˜10.14 for discussion on how the setting of the theorem is equivalent to the setting of Vershik-Kerov sequences.
@@
Afterwards, using Lemma˜10.7 gives that
@@
It is then is clear that $\zeta\triangleq\{J_{a(N)}^{A^{N-1}(\theta)}(x_{1},\ldots,x_{r},0,\ldots,0)\}_{N\geq r}$ is uniformly bounded over compact subsets of $\{z\in\mathbb{C}^{r}:|\text{Re}(z_{i})|<\frac{\theta}{rM}\,\forall i\in[r]\}$. Following this, we can use Corollary˜9.2 to deduce that the limit of the coefficient of $r(x)$ in $\zeta$ equals its coefficient in $\exp\left(\sum_{k\geq 1}\frac{m_{k}p_{(k)}\theta^{-k+1}}{k}\right)$; the application of Corollary˜9.2 is the difference between this proof and that of [rank_infinity]. We can then apply part (B) of Lemma˜10.8 to conclude the proof.
(B) and (C): We follow the same method as the proof of (A). For (B), we use part (B) of Lemma˜10.7 and Corollary˜9.6 and for (C), we use part (C) of Lemma˜10.7 and Corollary˜9.11. ∎
@@
From [assiotis]*Proposition 2.3, the condition that $\lim_{N\rightarrow\infty}\frac{\sum_{i=1}^{N}a(N)_{i}^{d}}{N^{d}}=m_{d}$ for all $d\geq 1$ implies that $\{a(N)\}_{N\geq 1}$ is a Vershik-Kerov sequence after it is reordered. Then, from [rank_infinity], we can set $M=\alpha+\epsilon$ in Theorem˜10.13 for any $\epsilon>0$, where $\alpha=\lim_{N\rightarrow\infty}\max_{i\in[N]}\left|\frac{a(N)_{i}}{N}\right|$. This will imply uniform convergence over compact subsets of $\{z\in\mathbb{C}^{r}:|\text{Re}(z_{i})|<\frac{\theta}{r\alpha}\}$, which is the version of the result that appears in [assiotis, rank_infinity].
@@ 10.5. Bessel generating functions for exponentially decaying probability measures
Lemma 10.16 ((A) is [limits_prob_measures]*Lemma 1.4 and (E) is [matrix]*Proposition 2.11).
@@
    The Bessel generating function $G_{\mu}^{\mathcal{R}(\theta)}(x)$ converges and is holomorphic over the closed ball of radius $R$ centered at the origin.
@@
    There exist unique constants $c_{\lambda}^{\mathcal{R}(\theta)}(\mu)\in\mathbb{C}$ for $\lambda\in\Gamma_{N}$ such that
    $G_{\mu}^{\mathcal{R}(\theta)}(x)=\exp\left(\sum_{\lambda\in\Gamma_{N}}c_{\lambda}^{\mathcal{R}(\theta)}(\mu)p_{\lambda}\right)$
@@
    There exist unique constants $m_{\epsilon}^{\mathcal{R}(\theta)}(\mu)\in\mathbb{C}$ for $\epsilon\in\mathcal{P}_{N}$ such that
    $G_{\mu}^{\mathcal{R}(\theta)}(x)=\exp\left(\sum_{\epsilon\in\mathcal{P}_{N}}m_{\epsilon}^{\mathcal{R}(\theta)}(\mu)M_{\epsilon}\right)$
    in a neighborhood of the origin.
  4. (D)
@@
For (B) and (C), we note that $G_{\mu}^{\mathcal{R}(\theta)}(x)$ is holomorphic in a neighborhood of the origin and $G_{\mu}^{\mathcal{R}(\theta)}(0)=1$.
(D): We include a detailed proof of this statement for completeness. We prove this result for the nonsymmetric eigenfunction $E_{a}^{\mathcal{R}(\theta)}(x)$. It is straightforward to deduce the result for the Bessel function afterwards using the expression relating the two functions in Theorem˜2.8. Note that some of the ideas of this proof have previously appeared in [limits_prob_measures]*Remark 4.6.
@@
converges because $|E_{a}^{\mathcal{R}(\theta)}(x)|\leq\sqrt{|H(\mathcal{R})|}\exp(\left\lVert a\right\rVert_{2}\left\lVert x\right\rVert_{2})$ and $\mu$ exponentially decays at rate $R$. Furthermore, we can show that the integral is holomorphic over $\overline{B}_{R}(0)$ by following [matrix]*Proof of Lemma 2.9. Note that we have just proved the nonsymmetric version of (A), which can be used to prove (A).
@@
(E): This is straightforward to deduce from (D). ∎
@@
The papers [matrix, limits_prob_measures] only consider the $A^{N}$ root system. However, (A) and (E) are generalizable to any finite root system after using the method discussed in [limits_prob_measures] that involves applying the results of [dunklbound]. Furthermore, observe that (A) implies (B) and (C). However, we require the exponentially decaying condition to prove (D) and (E).
Using Lemma˜10.16 and part (A) of Theorem˜1.1, we are able to deduce the following corollary for the $A^{N-1}$ root system. It is also straightforward to deduce the analogous results for the $BC^{N}$ and $D^{N}$ root systems. The corollary generalizes the results of [limits_prob_measures].
@@
Define $G_{\mu}^{A^{N-1}(\theta)}$, $c_{\lambda}^{A^{N-1}(\theta)}(\mu_{N})\in\mathbb{C}$ for $\lambda\in\Gamma_{N}$, and $m_{\epsilon}^{A^{N-1}(\theta)}(\mu_{N})\in\mathbb{C}$ for $\epsilon\in\mathcal{P}_{N}$ as they are defined in Lemma˜10.16. Then, the following are equivalent.
@@
    For all $\lambda\in\Gamma$, $\lim_{N\rightarrow\infty}\frac{c_{\lambda}^{A^{N-1}(\theta)}(\mu_{N})}{\theta N}=c_{\lambda}\in\mathbb{C}$ if $\ell(\lambda)=1$ and $\lim_{N\rightarrow\infty}\frac{c_{\lambda}^{A^{N-1}(\theta)}(\mu_{N})}{(\theta N)^{\ell(\lambda)}}=0$ if $\ell(\lambda)\geq 2$.
@@
    For all $\lambda\in\Gamma$, $\lim_{N\rightarrow\infty}\frac{\sum_{\epsilon\in\mathcal{P}_{N}}c_{\epsilon\lambda}m_{\epsilon}^{A^{N-1}(\theta)}(\mu_{N})}{\theta N}=c_{\lambda}\in\mathbb{C}$ if $\ell(\lambda)=1$ and $\lim_{N\rightarrow\infty}\\ \frac{\sum_{\epsilon\in\mathcal{P}_{N}}c_{\epsilon\lambda}m_{\epsilon}^{A^{N-1}(\theta)}(\mu_{N})}{(\theta N)^{\ell(\lambda)}}=0$ if $\ell(\lambda)\geq 2$.
  3. (a”)
    For all $\lambda\in\Gamma$, $\lim_{N\rightarrow\infty}\frac{\sum_{\epsilon\in\mathcal{P}_{N}}c_{\epsilon\lambda}(\epsilon!)^{-1}[1]\partial^{\epsilon}G_{\mu_{N}}^{A^{N-1}(\theta)}}{\theta N}=c_{\lambda}\in\mathbb{C}$ if $\ell(\lambda)=1$ and $\lim_{N\rightarrow\infty}\frac{\sum_{\epsilon\in\mathcal{P}_{N}}c_{\epsilon\lambda}(\epsilon!)^{-1}[1]\partial^{\epsilon}G_{\mu_{N}}^{A^{N-1}(\theta)}}{(\theta N)^{\ell(\lambda)}}=0$ if $\ell(\lambda)\geq 2$.
@@
Condition (b) of Corollary˜10.18 has been studied in previous works such as [llnclt, matrix, limits_prob_measures, rectangularmatrix, jackgenfunc]. When the condition is satisfied, the sequence of probability measures satisfies a law of large numbers.
@@
    Suppose $\mu_{a}^{N}$ and $\mu_{b}^{N}$ are exponentially decaying Borel probability measures over $\mathbb{R}^{N}$ for $N\geq 1$ such that:
@@
      The measure $\operatorname{\mathbb{E}}_{a\sim\mu_{a}^{N}}\left[\sum_{i=1}^{N}\frac{1}{N}\delta\left(\frac{a_{i}}{\theta N}\right)\right]$ converges to the compactly supported Borel probability measure $\mu_{a}$ over $\mathbb{R}$ in terms of moments.
@@
      The measure $\operatorname{\mathbb{E}}_{b\sim\mu_{b}^{N}}\left[\sum_{i=1}^{N}\frac{1}{N}\delta\left(\frac{b_{i}}{\theta N}\right)\right]$ converges to the compactly supported Borel probability measure $\mu_{b}$ over $\mathbb{R}$ in terms of moments.
@@
    Suppose $\mu_{a}^{N}$ and $\mu_{b}^{N}$ are compactly supported Borel probability measures over $\mathbb{R}^{N}$ for $N\geq 1$ and $\mu_{a}$ and $\mu_{b}$ are compactly supported Borel probability measures over $\mathbb{R}$ such that conditions 1 and 2 from (A) are true. For $N\geq 1$, suppose $\mu^{N}$ is a Borel probability measure such that $G_{\mu^{N}}^{A^{N-1}(\theta)}=G_{\mu_{a}^{N}}^{A^{N-1}(\theta)}G_{\mu_{b}^{N}}^{A^{N-1}(\theta)}$ in a neighborhood of the origin. Then, the measure $\mathbb{E}_{a\sim\mu^{N}}\left[\sum_{i=1}^{N}\frac{1}{N}\delta\left(\frac{a_{i}}{\theta N}\right)\right]$ converges to the free convolution of $\mu_{a}$ and $\mu_{b}$ in distribution and in terms of moments.

### v4

@@ 10.4. The uniform convergence of Bessel functions for Vershik-Kerov
In this subsection, we reprove the results of the papers [assiotis, rank_infinity] and extend these results to the $D^{N}$ root system. The setting that we consider is equivalent to the setting of Vershik-Kerov sequences that the papers consider, see Remark˜10.13. First, we prove the analogues of Theorems˜6.10, 7.4 and 8.21.
@@
Suppose $F_{N}(x_{1},\ldots,x_{N})=\exp\left(\sum_{\lambda\in\Gamma_{\text{even};\,N}}c_{\lambda}(N)p_{\lambda}\right)+ed_{0}(N)+e\exp\\ \left(\sum_{\lambda\in\Gamma_{\text{even};\,N}}d_{\lambda}(N)p_{\lambda}\right)$ for all $N\geq 1$. Assume that $\theta\in\mathbb{C}$ is fixed. Consider the following statements.
@@
    It is the case that $\lim_{N\rightarrow\infty}d_{0}(N)=d_{0}\in\mathbb{C}$ and for all $\lambda\in\Gamma_{\text{even}}$, $\lim_{N\rightarrow\infty}d_{\lambda}(N)=d_{\lambda}\in\mathbb{C}$.
@@
    It is the case that
    $\lim_{N\rightarrow\infty}\frac{1}{\prod_{j=1}^{N}(1+2(j-1)\theta)}[e,F_{N}]_{D^{N}(\theta)}=d_{0}+1$
    and for all $\nu\in\Gamma_{\text{even}}$
@@
Theorem˜10.12 is an example of an application of the results of this paper. Part (A) of the theorem appears in [assiotis] while parts (A) and (B) appear in [rank_infinity]. Part (C) has not appeared previously, although it is straightforward to deduce. See Remark˜10.13 for discussion on how the setting of the theorem is equivalent to the setting of Vershik-Kerov sequences.
@@
Afterwards, using Lemma˜10.6 gives that
@@
It is then is clear that $\zeta\triangleq\{J_{a(N)}^{A^{N-1}(\theta)}(x_{1},\ldots,x_{r},0,\ldots,0)\}_{N\geq r}$ is uniformly bounded over compact subsets of $\{z\in\mathbb{C}^{r}:|\text{Re}(z_{i})|<\frac{\theta}{rM}\,\forall i\in[r]\}$. Following this, we can use Corollary˜9.2 to deduce that the limit of the coefficient of $r(x)$ in $\zeta$ equals its coefficient in $\exp\left(\sum_{k\geq 1}\frac{m_{k}p_{(k)}\theta^{-k+1}}{k}\right)$; the application of Corollary˜9.2 is the difference between this proof and that of [rank_infinity]. We can then apply part (B) of Lemma˜10.7 to conclude the proof.
(B) and (C): We follow the same method as the proof of (A). For (B), we use part (B) of Lemma˜10.6 and Corollary˜9.6 and for (C), we use part (C) of Lemma˜10.6 and Corollary˜9.11. ∎
@@
From [assiotis]*Proposition 2.3, the condition that $\lim_{N\rightarrow\infty}\frac{\sum_{i=1}^{N}a(N)_{i}^{d}}{N^{d}}=m_{d}$ for all $d\geq 1$ implies that $\{a(N)\}_{N\geq 1}$ is a Vershik-Kerov sequence after it is reordered. Then, from [rank_infinity], we can set $M=\alpha+\epsilon$ in Theorem˜10.12 for any $\epsilon>0$, where $\alpha=\lim_{N\rightarrow\infty}\max_{i\in[N]}\left|\frac{a(N)_{i}}{N}\right|$. This will imply uniform convergence over compact subsets of $\{z\in\mathbb{C}^{r}:|\text{Re}(z_{i})|<\frac{\theta}{r\alpha}\}$, which is the version of the result that appears in [assiotis, rank_infinity].
@@ 10.5. Bessel generating functions for exponentially decaying probability measures
Lemma 10.15 ((A) is [limits_prob_measures]*Lemma 1.4 and (D) is [matrix]*Proposition 2.11).
@@
    The Bessel generating function $G_{\mu}^{\mathcal{R}(\theta)}(x)$ converges over the closed ball of radius $R$ centered at the origin and is holomorphic over the interior of this domain.
@@
    Let $\{1\}\cup\mathcal{B}$ be a basis of $\mathbb{C}^{H(\mathcal{R})}[x_{1},\ldots,x_{N}]$. There exist unique constants $c_{b}^{\mathcal{R}(\theta)}(\mu)\in\mathbb{C}$ for $b\in\mathcal{B}$ such that
    $G_{\mu}^{\mathcal{R}(\theta)}(x)=\exp\left(\sum_{b\in\mathcal{B}}c_{b}^{\mathcal{R}(\theta)}(\mu)b\right)$
@@
For (B), we note that $G_{\mu}^{\mathcal{R}(\theta)}(x)$ is holomorphic in a neighborhood of the origin and $G_{\mu}^{\mathcal{R}(\theta)}(0)=1$.
(C): We include a detailed proof of this statement for completeness. We prove this result for the nonsymmetric eigenfunction $E_{a}^{\mathcal{R}(\theta)}(x)$. It is straightforward to deduce the result for the Bessel function afterwards using the expression relating the two functions in Theorem˜2.8. Note that some of the ideas of this proof have previously appeared in [limits_prob_measures]*Remark 4.6.
@@
converges because $|E_{a}^{\mathcal{R}(\theta)}(x)|\leq\sqrt{|H(\mathcal{R})|}\exp(\left\lVert a\right\rVert_{2}\left\lVert x\right\rVert_{2})$ and $\mu$ exponentially decays at rate $R$. Furthermore, we can show that the integral is holomorphic over $B_{R}(0)$ by following [matrix]*Proof of Lemma 2.9. Note that we have just proved the nonsymmetric version of (A), which can be used to prove (A).
@@
(D): This is straightforward to deduce from (C). ∎
@@
The papers [matrix, limits_prob_measures] only consider the $A^{N}$ root system. However, (A) and (D) are generalizable to any finite root system after using the method discussed in [limits_prob_measures] that involves applying the results of [dunklbound]. Furthermore, observe that (A) implies (B). However, we require the exponentially decaying condition to prove (C) and (D).
Using Lemma˜10.15 and part (A) of Theorem˜1.1, we are able to deduce the following corollary for the $A^{N-1}$ root system. It is also straightforward to deduce the analogous results for the $BC^{N}$ and $D^{N}$ root systems. The corollary generalizes the results of [limits_prob_measures].
@@
For $\mathcal{B}=\{p_{\lambda}:\lambda\in\Gamma_{N}\}$, define the coefficients $c_{p_{\lambda}}^{A^{N-1}(\theta)}(\mu_{N})\in\mathbb{C}$ for $\lambda\in\Gamma_{N}$ and for $\mathcal{B}=\{M_{\epsilon}:\epsilon\in\mathcal{P}_{N}\}$, define the coefficients $c_{M_{\epsilon}}^{A^{N-1}(\theta)}(\mu_{N})\in\mathbb{C}$ for $\epsilon\in\mathcal{P}_{N}$ as they are defined in part (B) of Lemma˜10.15. Then, the following are equivalent.
@@
    For all $\lambda\in\Gamma$, $\lim_{N\rightarrow\infty}\frac{c_{p_{\lambda}}^{A^{N-1}(\theta)}(\mu_{N})}{\theta N}=c_{\lambda}\in\mathbb{C}$ if $\ell(\lambda)=1$ and $\lim_{N\rightarrow\infty}\frac{c_{p_{\lambda}}^{A^{N-1}(\theta)}(\mu_{N})}{(\theta N)^{\ell(\lambda)}}=0$ if $\ell(\lambda)\geq 2$.
@@
    For all $\lambda\in\Gamma$, $\lim_{N\rightarrow\infty}\frac{\sum_{\epsilon\in\mathcal{P}_{N}}c_{\epsilon\lambda}c_{M_{\epsilon}}^{A^{N-1}(\theta)}(\mu_{N})}{\theta N}=c_{\lambda}\in\mathbb{C}$ if $\ell(\lambda)=1$ and $\lim_{N\rightarrow\infty}\\ \frac{\sum_{\epsilon\in\mathcal{P}_{N}}c_{\epsilon\lambda}c_{M_{\epsilon}}^{A^{N-1}(\theta)}(\mu_{N})}{(\theta N)^{\ell(\lambda)}}=0$ if $\ell(\lambda)\geq 2$.
@@
Condition (b) of Corollary˜10.17 has been studied in previous works such as [llnclt, matrix, limits_prob_measures, rectangularmatrix, jackgenfunc]. When the condition is satisfied, the sequence of probability measures satisfies a law of large numbers. Next, we state the analogous result for the $BC^{N}$ root system.
Corollary 10.18.
Suppose $\mu_{N}$ is an exponentially decaying Borel probability measure over $\mathbb{C}^{N}$ for all $N\geq 1$. Assume that $\theta_{0},\theta_{1}\in\mathbb{C}$ have nonnegative real part for all $N\geq 1$, $\lim_{N\rightarrow\infty}|\theta_{0}N|=\infty$, and $\lim_{N\rightarrow\infty}\frac{\theta_{1}}{\theta_{0}N}=c\in\mathbb{C}$.
For $\mathcal{B}=\{p_{\lambda}:\lambda\in\Gamma_{\text{even};\,N}\}$, define the coefficients $c_{p_{\lambda}}^{BC^{N}(\theta_{0},\theta_{1})}(\mu_{N})\in\mathbb{C}$ for $\lambda\in\Gamma_{\text{even};\,N}$ and for $\mathcal{B}=\{M_{\epsilon}:\epsilon\in\mathcal{P}_{\text{even};\,N}\}$, define the coefficients $c_{M_{\epsilon}}^{BC^{N}(\theta_{0},\theta_{1})}(\mu_{N})\in\mathbb{C}$ for $\epsilon\in\mathcal{P}_{\text{even};\,N}$ as they are defined in part (B) of Lemma˜10.15. Then, the following are equivalent.
  1. (a)
    For all $\lambda\in\Gamma_{\text{even}}$, $\lim_{N\rightarrow\infty}\frac{c_{p_{\lambda}}^{BC^{N}(\theta_{0},\theta_{1})}(\mu_{N})}{\theta N}=c_{\lambda}\in\mathbb{C}$ if $\ell(\lambda)=1$ and $\lim_{N\rightarrow\infty}\\ \frac{c_{p_{\lambda}}^{BC^{N}(\theta_{0},\theta_{1})}(\mu_{N})}{(\theta N)^{\ell(\lambda)}}=0$ if $\ell(\lambda)\geq 2$.
  2. (a’)
    For all $\lambda\in\Gamma_{\text{even}}$, $\lim_{N\rightarrow\infty}\frac{\sum_{\epsilon\in\mathcal{P}_{\text{even};\,N}}c_{\epsilon\lambda}c_{M_{\epsilon}}^{BC^{N}(\theta_{0},\theta_{1})}(\mu_{N})}{\theta N}=c_{\lambda}\in\mathbb{C}$ if $\ell(\lambda)=1$ and $\lim_{N\rightarrow\infty}\frac{\sum_{\epsilon\in\mathcal{P}_{\text{even};\,N}}c_{\epsilon\lambda}c_{M_{\epsilon}}^{BC^{N}(\theta_{0},\theta_{1})}(\mu_{N})}{(\theta N)^{\ell(\lambda)}}=0$ if $\ell(\lambda)\geq 2$.
  3. (b)
    For all $\nu\in\Gamma_{\text{even}}$,
    $\lim_{N\rightarrow\infty}\frac{1}{N^{\ell(\nu)}}\operatorname{\mathbb{E}}_{a\sim\mu_{N}}\left[p_{\nu}\left(\frac{a}{\theta_{0}N}\right)\right]=\prod_{i=1}^{\ell(\nu)}\sum_{\pi\in NC^{\text{even}}(\nu_{i})}(1+c)^{o(\pi)}\prod_{B\in\pi}2^{|B|-1}|B|c_{(|B|)}.$
For the type D analogue of Corollary˜10.17, we adjust the notation used to match the notation of part (C) of Theorem˜1.1. Therefore, we first prove the following lemma, which is analogous to part (B) of Lemma˜10.15.
Lemma 10.19.
Suppose the Borel probability measure $\mu$ over $\mathbb{C}^{N}$ is exponentially decaying at rate $R>0$. Furthermore, assume that $\theta\in\mathbb{C}$ has nonnegative real part.
Let $\{1\}\cup\{\mathcal{B}\}$ be a basis of $\mathbb{C}^{H(BC^{N})}[x_{1},\ldots,x_{N}]$. There exist unique constants $d_{0}^{D^{N}(\theta)}(\mu)\\$ $\in\mathbb{C}$ and $c_{b}^{D^{N}(\theta)}(\mu),d_{b}^{D^{N}(\theta)}(\mu)\in\mathbb{C}$ for $b\in\mathcal{B}$ such that
$G_{\mu}^{D^{N}(\theta)}=\exp\left(\sum_{b\in\mathcal{B}}c_{b}^{\mathcal{R}(\theta)}(\mu)b\right)+ed_{0}^{D^{N}(\theta)}(\mu)+e\exp\left(\sum_{b\in\mathcal{B}}c_{b}^{\mathcal{R}(\theta)}(\mu)b\right)$
over a neighborhood of the origin.
Proof.
We have that $\frac{1}{2}(G_{\mu}^{D^{N}(\theta)}(x)+G_{\mu}^{D^{N}(\theta)}(-x))$ is even and evaluates to one at the origin, from which the uniqueness of $c_{b}^{D^{N}(\theta)}(\mu)$ for $b\in\mathcal{B}$ follows. Furthermore, we have that $\frac{1}{2}(G_{\mu}^{D^{N}(\theta)}(x)-G_{\mu}^{D^{N}(\theta)}(-x))$ is odd and $\frac{G_{\mu}^{D^{N}(\theta)}(x)-G_{\mu}^{D^{N}(\theta)}(-x)}{2e(x)}$ is even. It is clear that $d_{0}^{D^{N}(\theta)}(\mu)=\frac{G_{\mu}^{D^{N}(\theta)}(x)-G_{\mu}^{D^{N}(\theta)}(-x)}{2e(x)}-1$. Next, we have that $\frac{G_{\mu}^{D^{N}(\theta)}(x)-G_{\mu}^{D^{N}(\theta)}(-x)}{2e(x)}-d_{0}^{D^{N}(\theta)}(\mu)$ evaluates to one at the origin, from which the uniqueness of $d_{b}^{D^{N}(\theta)}(\mu)$ for $b\in\mathcal{B}$ follows. ∎
Corollary 10.20.
Define the coefficient $d_{0}^{D^{N}(\theta)}(\mu)$, the coefficients $c_{p_{\lambda}}^{D^{N}(\theta)}(\mu_{N}),d_{p_{\lambda}}^{D^{N}(\theta)}(\mu_{N})\in\mathbb{C}$ for $\lambda\in\Gamma_{\text{even};\,N}$ for $\mathcal{B}=\{p_{\lambda}:\lambda\in\Gamma_{\text{even};\,N}\}$, and the coefficients $c_{M_{\epsilon}}^{D^{N}(\theta)}(\mu_{N}),d_{M_{\epsilon}}^{D^{N}(\theta)}(\mu_{N})\in\mathbb{C}$ for $\epsilon\in\mathcal{P}_{\text{even};\,N}$ for $\mathcal{B}=\{M_{\epsilon}:\epsilon\in\mathcal{P}_{\text{even};\,N}\}$ as they are defined in Lemma˜10.19. Consider the following statements.
  1. (a)
    For all $\lambda\in\Gamma_{\text{even}}$, $\lim_{N\rightarrow\infty}\frac{c_{p_{\lambda}}^{D^{N}(\theta)}(\mu_{N})}{\theta N}=c_{\lambda}\in\mathbb{C}$ if $\ell(\lambda)=1$ and $\lim_{N\rightarrow\infty}\frac{c_{p_{\lambda}}^{D^{N}(\theta)}(\mu_{N})}{(\theta N)^{\ell(\lambda)}}\\ =0$ if $\ell(\lambda)\geq 2$.
  2. (a’)
    For all $\lambda\in\Gamma_{\text{even}}$, $\lim_{N\rightarrow\infty}\frac{\sum_{\epsilon\in\mathcal{P}_{\text{even};\,N}}c_{\epsilon\lambda}c_{M_{\epsilon}}^{D^{N}(\theta)}(\mu_{N})}{\theta N}=c_{\lambda}\in\mathbb{C}$ if $\ell(\lambda)=1$ and $\lim_{N\rightarrow\infty}\frac{\sum_{\epsilon\in\mathcal{P}_{\text{even};\,N}}c_{\epsilon\lambda}c_{M_{\epsilon}}^{D^{N}(\theta)}(\mu_{N})}{(\theta N)^{\ell(\lambda)}}=0$ if $\ell(\lambda)\geq 2$.
  3. (b)
    It is the case that $\lim_{N\rightarrow\infty}d_{0}(N)=d_{0}\in\mathbb{C}$ and for all $\lambda\in\Gamma_{\text{even};\,N}$, $\lim_{N\rightarrow\infty}\\ \frac{d_{p_{\lambda}}^{D^{N}(\theta)}(\mu_{N})}{\theta N}=d_{\lambda}\in\mathbb{C}$ if $\ell(\lambda)=1$ and $\lim_{N\rightarrow\infty}\frac{d_{p_{\lambda}}^{D^{N}(\theta)}(\mu_{N})}{(\theta N)^{\ell(\lambda)}}=0$ if $\ell(\lambda)\geq 2$.
  4. (b’)
    It is the case that $\lim_{N\rightarrow\infty}d_{0}(N)=d_{0}\in\mathbb{C}$ and for all $\lambda\in\Gamma_{\text{even};\,N}$, $\lim_{N\rightarrow\infty}\\ \frac{\sum_{\epsilon\in\mathcal{P}_{\text{even};\,N}}c_{\epsilon\lambda}d_{M_{\epsilon}}^{D^{N}(\theta)}(\mu_{N})}{\theta N}=d_{\lambda}\in\mathbb{C}$ if $\ell(\lambda)=1$ and $\lim_{N\rightarrow\infty}\frac{\sum_{\epsilon\in\mathcal{P}_{\text{even};\,N}}c_{\epsilon\lambda}d_{M_{\epsilon}}^{D^{N}(\theta)}(\mu_{N})}{(\theta N)^{\ell(\lambda)}}=0$ if $\ell(\lambda)\geq 2$.
  5. (c)
    For all $\nu\in\Gamma_{\text{even}}$,
    $\lim_{N\rightarrow\infty}\frac{1}{N^{\ell(\nu)}}\operatorname{\mathbb{E}}_{a\sim\mu_{N}}\left[p_{\nu}\left(\frac{a}{\theta N}\right)\right]=\prod_{i=1}^{\ell(\nu)}\sum_{\pi\in NC^{\text{even}}(\nu_{i})}\prod_{B\in\pi}2^{|B|-1}|B|c_{(|B|)}.$
  6. (d)
    It is the case that
    $\lim_{N\rightarrow\infty}\frac{\operatorname{\mathbb{E}}_{a\sim\mu_{N}}\left[e(a)\right]}{\prod_{j=1}^{N}(1+2(j-1)\theta)}=d_{0}+1$
    and for all $\nu\in\Gamma_{\text{even}}$,
    $\lim_{N\rightarrow\infty}\frac{\operatorname{\mathbb{E}}_{a\sim\mu_{N}}\left[e(a)p_{\nu}\left(\frac{a}{\theta N}\right)\right]}{N^{\ell(\nu)}\prod_{j=1}^{N}(1+2(j-1)\theta)}=\prod_{i=1}^{\ell(\nu)}\sum_{\pi\in NC^{\text{even}}(\nu_{i})}\prod_{B\in\pi}2^{|B|-1}|B|d_{(|B|)}.$
@@
[…]
@@
      The measure $\operatorname{\mathbb{E}}_{a\sim\mu_{a}^{N}}\left[\sum_{i=1}^{N}\frac{1}{N}\delta\left(\frac{a_{i}}{\theta N}\right)\right]$ converges to the compactly supported Borel probability measure $\mu_{a}$ over $\mathbb{C}$ in terms of moments.
@@
      The measure $\operatorname{\mathbb{E}}_{b\sim\mu_{b}^{N}}\left[\sum_{i=1}^{N}\frac{1}{N}\delta\left(\frac{b_{i}}{\theta N}\right)\right]$ converges to the compactly supported Borel probability measure $\mu_{b}$ over $\mathbb{C}$ in terms of moments.
@@
    Suppose $\theta\in\mathbb{C}$ is positive. Suppose $\mu_{a}^{N}$ and $\mu_{b}^{N}$ are compactly supported Borel probability measures over $\mathbb{R}^{N}$ for $N\geq 1$ and $\mu_{a}$ and $\mu_{b}$ are compactly supported Borel probability measures over $\mathbb{R}$ such that conditions 1 and 2 from (A) are true. For $N\geq 1$, suppose $\mu^{N}$ is a Borel probability measure over $\mathbb{R}^{N}$ such that $G_{\mu^{N}}^{A^{N-1}(\theta)}=G_{\mu_{a}^{N}}^{A^{N-1}(\theta)}G_{\mu_{b}^{N}}^{A^{N-1}(\theta)}$ in a neighborhood of the origin. Then, the measure $\mathbb{E}_{a\sim\mu^{N}}\left[\sum_{i=1}^{N}\frac{1}{N}\delta\left(\frac{a_{i}}{\theta N}\right)\right]$ converges to the free convolution of $\mu_{a}$ and $\mu_{b}$ in distribution and in terms of moments.